AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Thu May 29 18:04:47 2025
Author
*
:
Subject
*
:
<p>Oven temperature increased, current up to 3.5E7 pps.</p> <p>Found beam-target overlap setting horizontal bump in RO9 to <strong>-2.5mm.</strong></p> <p>Target T1 30.5K, density increasing to 2.6E11 at/cm2.</p> <p><strong>19:08 Run 32 start</strong></p> <p>Moving left detectors (79.69->34.66mm).</p> <p>Beam energy 1.85 MeV/u nominal</p> <p>Electron cooler actual values: voltage 1.051 kV, current 0.013215 uA, [voltage divider G35 1.0504 kV]</p> <p>Stopped run, beam was still in bunch mode.</p> <p><strong>19:15 Run 33 start</strong></p> <p>All checks Ok.</p> <p>Target at 3E11 at/cm2</p> <p>Attachment 1-2 with rates and temperatures</p> <p> </p> <p>20:20</p> <p>Target density 3.4E11 at/cm2. Beam around 2.2-3.6E7 pps.</p> <p>Checks ok! Spectra attached in att.3.</p> <p>21:25</p> <p>all good. Target density 4E11 at/cm^2</p> <p>22:24</p> <p>All checks are OK</p> <p>Noted some variability in the injection intensity. For 15 minutes about half of shots were below 1e7. Perhaps some variability due to end of lithium in oven or increased temperture. </p> <p>Beam calc with read values - Attachment 4</p> <p>01:30</p> <p>All checks OK. Injection intensity has stabilised ~2E7</p> <p>Target density is stable ~4e11 at/cm2</p> <p>02:44</p> <p>Suddenly see very high rates in DSSD4. Detector had tripped. Tried to re-bias but kept tripping.</p> <p>Set to 70V, leakage current is about 17uA but fluctuates quite wildly. Set to 50V and leakage current is about 6uA but again has wild fluctuations. Not sure of the cause, this detector is not moving and is far from the beam.</p> <p>Decided to turn detector off to try and protect it. We can re-bias and monitor in the day tomorrow when we do not have beam on target.</p> <p>This detector has tripped previously but was able to be re-biased at the time. See attachments 5 and 6 for the detector leakage currents.</p> <p><strong>New Run started R34. </strong></p> <p>Same conditions as before. DSSD4 turned off. To stop crap writing to disk I have set the corresponding FEEs to the MEC.</p> <p>In principle we do not need this detector now as it is not moving to see the reaction.</p> <p>Note I also observed in the system messages we semi regularly get a DHCP request from aida11. It seems to be functioning fine as I see the pulser signal in the spectrum browser. May be worth following up on at a later date. See attachment 7</p> <p>All other checks OK</p> <p>Data rate is now ~140 kB/sec</p> <p>03:15</p> <p>Tried to bias DSSD4 again. Leakage currents appear fine again.</p> <p>Left for 5 minutes to gauge stability -> seems fine.</p> <p>03:22</p> <p>Have returned aida13, 14, 15, 16 back to LEC. Maintaing the same run (34)</p> <p>Scratch that. 03:27 detector tripped again. Will just leave it off tonight. Turned those FEEs back to MEC.</p> <p>06:00</p> <p>I have quickly looked at Run 33. We clearly see the 7Be again in both detectors. Locus on detector much smaller than at higher energies. </p> <p>All checks OK</p> <p>07:07</p> <p>Run stopped, switching to deuterium.</p> <p>Detectors moved out</p> <p>Shutter for gas target off.</p> <p>Moved to next run -> no storage mode.</p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Resubmit as new entry
Attachment 1:
Screenshot_from_2025-05-29_19-16-09.png
Original size: 1914x466
Attachment 2:
Screenshot_from_2025-05-29_19-12-46.png
Original size: 1913x544
Attachment 3:
Screenshot_from_2025-05-29_20-23-50.png
Original size: 1699x911
Attachment 4:
beamCalc.png
Original size: 516x853
Attachment 5:
Screenshot_from_2025-05-30_02-37-55.png
Original size: 725x462
Attachment 6:
Screenshot_from_2025-05-30_02-51-21.png
Original size: 1920x977
Attachment 7:
Screenshot_from_2025-05-30_02-38-10.png
Original size: 1742x462
Attachment 8:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.3-7933898